3.f(x)=2-3x solve to f(-3),f(-1) and f(4)
Answer by ikleyn(52781)

.

"Solve to f(-3)"  means  "substitute the value -3 instead of x into the formula and calculate"


    f(-3) = 2 - 3*(-3) = 2 + 3*3 = 2 + 9 = 11.    ANSWER

------------

First example is done for you.

Do the rest on your own in the same way.


Generally speaking, these questions are NOT INTENDED to go to the forum and to ask for help from outside.


        Making calculations is the students' honor duty and a privilege (!)